OpenStudy (anonymous):

what is the vertex of the following problem f(x)=2x^2-12x+8

OpenStudy (across):

\[f(x)=2x^2-12x+8\]\[f'(x)=4x-12=0\]\[4x=12\]\[x=3\]

OpenStudy (across):

The vertex is then at\[f(3)=2(3)^2-12(3)+8\]

OpenStudy (across):

Write it as an ordered pair:\[(3,f(3))\]

OpenStudy (anonymous):

so the vertex is then (3,8) right

OpenStudy (anonymous):

alright it is (3,f3) i dont understand

OpenStudy (across):

The vertex should be at\[(3,-10)\](Evaluate the above function).